# Double-Signed Fragmented DNSSEC for Countering Quantum Threat

### v1

@@ Abstract.
DNSSEC, a DNS security extension, is essential to accurately translating domain names to IP addresses. Digital signatures provide the foundation for this reliable translation, however, the evolution of Quantum Computers has made traditional digital signatures vulnerable. In light of this, NIST has recently selected potential post-quantum digital signatures that can operate on conventional computers and resist attacks made with Quantum Computers. Since these post-quantum digital signatures are still in their early stages of development, replacing pre-quantum digital signature schemes in DNSSEC with post-quantum candidates is risky until the post-quantum candidates have undergone a thorough security analysis. Given this, herein, we investigate the viability of employing Double-Signatures in DNSSEC, combining a post-quantum digital signature and a classic one. The rationale is that double-signatures will offer protection against quantum threats on conventional signature schemes as well as unknown non-quantum attacks on post-quantum signature schemes, hence even if one fails the other provides security guarantees. However, the inclusion of two signatures in the DNSSEC response message doesn’t bode well with the maximum allowed size of DNSSEC responses (i.e., 1232B, a limitation enforced by MTU of physical links). To counter this issue, we leverage a way to do application-layer fragmentation of DNSSEC responses with two signatures. We implement our solution on top of OQS-BIND and through experiments show that the addition of two signatures in DNSSEC and application-layer fragmentation of all relevant resource records and their reassembly does not have any substantial impact on the efficiency of the resolution process and thus is suitable for the interim period at least until the quantum computers are fully realized.
@@
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: Conference; Oct, 2024; Australiaisbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
The Domain Name System (DNS) is responsible for translating human-readable domain names into machine-understandable IP addresses. As can be imagined, any vulnerability in this system can potentially make such translation precarious, leading users to malicious servers instead of their intended destinations. The Security Extension of DNS - i.e., Domain Name System Security Extensions (DNSSEC) helps establish that the messages received by a client are from an authorized DNS server and have not been altered in transit. DNSSEC primarily leverages digital signatures for establishing the aforementioned properties of authenticity and integrity. However, the underlying security assumptions of conventional digital signature algorithms - i.e., the computational difficulty of integer factorization and discrete logarithms will not hold when an attacker has access to a Cryptographically Relevant Quantum Computer (CRQC) (Agency, 2021; Shor, 1994). To mitigate this problem, the National Institute of Standards and Technology (NIST) has recently shortlisted three post-quantum digital signatures - i.e., FALCON (Prest et al., 2020), CRYSTALS-DILITHIUM (Lyubashevsky et al., 2020), and SPHINCS+ (Hulsing et al., 2020). An important envisioned property of these post-quantum digital signature algorithms is that they can run on classical computers but also withstand the attacks conducted using CRQCs, thereby offering sufficient protection against the impending threat of CRQCs. For DNSSEC to continue offering the needed authenticity and integrity in the post-quantum era, it must be updated with post-quantum digital signatures. However, for successful transitioning to the post-quantum era, the Interim Period - i.e., from now until CRQCs are fully realized or until the security of post-quantum candidates is fully established, presents a particular challenge. Precisely, in preparation for the post-quantum era, simply switching pre-quantum cryptography (such as conventional digital signatures in DNSSEC) with post-quantum cryptography (such as the post-quantum candidates shortlisted by NIST) in the interim period may lower overall security (ENISA, 2022). The reason is that the shortlisted post-quantum candidates (both key encapsulation mechanisms and digital signatures) have yet not gone through a thorough cryptanalysis against both classical and quantum attacks. The shortlisted post-quantum candidates may be compromised even using conventional classical computers in the near future. Therefore, there is an inherent risk in simply switching from pre-quantum cryptography to post-quantum cryptography alone in that overall security may be lowered not only against CRQCs but also against today’s classical computers (see further details in Section 3).
Aligned with the aforementioned discussion and the recommendations from the European Union Agency For Cybersecurity (ENISA), we in this paper investigate the plausibility of Double-Signed DNSSEC by combining the pre-quantum and post-quantum digital signatures for the interim period. However, the size of signatures and public keys of post-quantum candidates shortlisted by NIST is very large (i.e., shortlisted post-quantum signatures are 11$\times$ to 122$\times$ and public keys are 14$\times$ to 20$\times$ of their pre-quantum counterparts). These sizes further increase when double signatures are incorporated into the DNSSEC resolution process to ensure authenticity and integrity. This increased size of signatures and public keys significantly impacts the DNSSEC resolution process (Kampanakis and Lepoint, 2023). Precisely, since UDP is the preferred transport protocol for DNSSEC because of it being lightweight, any response message that exceeds 1232B triggers IP-level fragmentation which is unreliable. The maximum size of DNSSEC response messages - i.e., 1232B is obtained as: (IPv6 MTU=1280)-(IPv6 Header=40)-(UPD Header=8). With double signatures, the sizes of DNSSEC response messages that carry different resource records for facilitating reliable address resolutions are much larger than the threshold size of 1232B. Therefore, these large DNSSEC messages necessitate IP fragmentation which is unreliable over UDP (i.e., some fragments may never reach the resolver) (Van Den Broek et al., 2014). Due to this issue, the standard DNSSEC offers a fallback mechanism via TCP for sending large response messages without resorting to fragmentation. Precisely, when the response size exceeds the threshold (i.e., the resolver’s advertised EDNS(0) buffer size), name servers send back a truncated response (i.e., by setting the TC flag), indicating the resolver to retry via TCP (and thus negatively contributing to the address resolution time). However, many name servers lack support for TCP in addition to high-resolution time with TCP due to it being not as lightweight as UDP (Van Den Broek et al., 2014; van Rijswijk-Deij et al., 2016; Mao et al., 2022).
@@
To circumvent the aforementioned issues, a potential alternative is to perform the fragmentation at the application layer (i.e., within the DNS protocol itself). While there exist ways to tackle the application layer fragmentation (see (Goertzen and Stebila, 2022; Rawat and Jhanwar, 2024)), none of them can accommodate double signatures and public keys. Therefore, to accomplish this, first, we constructed a Docker-based testbed on an Amazon EC2 instance using commercial grade DNS software - i.e., BIND9 (OQS-BIND, see full details on our testbed in Section 5). Second, we explored generating two signatures (pre-quantum and post-quantum) on the Name Server using the default BIND9 tools. Third, we devised a way to fragment the DNSSEC responses with two signatures and public keys on the name servers and accurately reassemble them on the resolver. We also modified the source code of the BIND9 resolver to enable the verification of both signatures before marking the response as authenticated. Finally, using our testbed, we conducted an empirical analysis to show that our implementation can handle two signatures in DNSSEC simultaneously and quantify the impact of double signatures on the DNSSEC resolution process. Our experiments show that double signatures have a negligible impact on the average resolution time of DNSSEC compared to only post-quantum digital signatures. Therefore, this analysis confirms the efficacy of double signatures in DNSSEC for the interim period until CRQCs are fully realized. Our main contributions are summarized as follows:
@@
  • We thoroughly investigated the source code for BIND9 and extensively modified the resolver component to enable the verifications of both pre-quantum and post-quantum digital signatures at the same time.
  • Through empirical analysis we show that the addition of double signatures in the DNSSEC resolution process did not affect the performance. Hence, it is recommended for the interim period until CRQCs are fully realized.
@@ 2. Domain Name System (DNS)
For an abstract-level understanding of the DNS translation process, imagine a client who queries the IP address of a domain name example.com. For this translation, the client will first send a query to a (caching) resolver which will handle all the subsequent translation process on the client’s behalf. Assuming that the resolver has not previously cached this answer (i.e., the IP address of example.com), it will query the root server for the name server of .com domain names. Once the resolver receives the response, it queries this name server for the name server(s) of example.com. Finally, once the resolver knows the corresponding name server, it queries it for the IP address of domain name example.com and returns it to the client.
As specified in RFC 1035, the generic DNS messages exchanged between the resolver and name servers have five sections (Mockapetris, 1987). Figure 1 shows an overview of a generic DNS message. The header section is always present and its structure is depicted in Figure 2. The Question section encapsulates the resolver’s query whose IP address needs to be resolved. This section has three main fields - i.e., QName (contains encoded query), QType (mentions the type of needed resource recorded in response), and QClass (contains the class of query, mostly specifies IN). As shown in Figure 1, the last three sections have a similar format - i.e., concatenation of various RRs. The RRs are actual data entries in the DNS database that provide information related to the queried domain name and help the actual translation of domain names to corresponding IP addresses. The top-level format of RRs is shown below in Figure 3. Going back to our earlier example of example.com, an RR with Name ‘example.com’ and TYPE ‘A’ will contain an IP address of 32 bits in the RDATA field, and with TYPE ‘NS’, RDATA field will contain the domain name of the authoritative server for example.com. The Answer section in a response message contains RRs that answer the query from the resolver, while the Authority section contains RRs that point to the authoritative server. The Additional section contains RRs that somehow relate to the query but may not essentially contain the answers to the query.
@@
Figure 3. RR’s Top-Level Format
@@
As explained above, the client initiates the translation process by querying the root server, then its children, and then its children’s children, and finally reaching the name server that provides the queried IP address. This hierarchical process facilitates an important element of DNSSEC - i.e., the chain of trust that helps establish trust in public keys needed for signature verification. For the chain of trust, each zone that is queried as part of the resolution process must have a digest of public KSK kept as the Delegation Signer resource record in its parent zone; otherwise, the public ZSK sent by the name server cannot be trusted. Since the root server(s) has no parent zone, it does not publish any delegation signer (DS) resource record. Therefore, for the entire chain of trust, the public KSK of the root server is obtained out-of-band so that the root server’s public KSK is pre-installed on most operating systems to eliminate the need for fetching and configuring the key(s). An overview of the chain of trust is shown in Figure 4.
@@
Since RRSIG and DNSKEY RRs are the root cause of the maximum message size issue related to the double signatures (see discussion in the introduction), below we present an overview of RRSIG and DNSKEY RRs (see Figures 5, 6). Their thorough understanding is pivotal for the simultaneous fragmentation of two signatures at the application layer. We refer the reader to RFC 4034 for a more detailed discussion on these relevant RRs (Arends et al., 2005).
Figure 5. RRSIG RR Wire Format
Figure 6. DNSKEY RR Wire Format
@@ 3.1. Quantum Threat
The assurances of source authenticity and data integrity in DNSSEC are underpinned by digital signatures. However, these conventional digital signatures used in DNSSEC rely on the complexity and difficulty of computational problems (e.g., integer factorization and discrete logarithms) which are difficult for attackers to solve using today’s classical computers (Nosouhi et al., 2023a, b). There exists strong evidence that these “computationally hard problems” will become solvable in “polynomial time” with the quantum computational model using the so-called CRQCs (Mosca, 2018; Agency, 2021). These CRQCs will facilitate the implementation of Shor’s Algorithm that will help solve the aforementioned computational problems in polynomial time (Shor, 1994), thereby rendering the conventional digital signatures insecure. For DNSSEC, it means that in the post-quantum era (i.e., when CRQCs will be fully realized), it would be difficult to establish the integrity and authenticity of all resource records exchanged as a part of the address resolution process. Therefore, in the post-quantum era, it will become impossible to trust the final IP address of a queried domain as the conventional digital signature will no longer offer the source authenticity and data integrity at an adequate level.
@@
Given the aforementioned impending issues with conventional digital signatures (and also with existing key encapsulation mechanisms), NIST is currently undertaking a Post Quantum Cryptography (PQC) standardization project. It aims to standardize digital signature algorithms and key encapsulation mechanisms that can run on today’s classical computers but can withstand the attacks conducted using the CRQCs (if they are realized in the future). In this pursuit, NIST has recently shortlisted three post-quantum digital signature algorithms, including FALCON (Prest et al., 2020), CRYSTALS-DILITHIUM (Lyubashevsky et al., 2020), and SPHINCS++ (Hulsing et al., 2020). To make DNSSEC (or in general any other internet protocol) quantum-safe, it must be updated with the aforementioned post-quantum digital signature algorithms and be investigated for any issues that demand protocol-level changes to DNSSEC for the post-quantum era. However, simply switching the conventional digital signatures with post-quantum algorithms may be precarious. The reason is that the post-quantum digital signature algorithms as shortlisted by NIST are currently in their infancy and thus have been studied and analyzed for a shorter period (as compared with their classical counterparts). First, there could be flaws in the algorithm design itself that make them vulnerable to cryptanalytic attacks. Second, there could be side-channels in the different implementations that might leak keys via timing attacks, power analysis attacks, and other similar attacks. Third, there could be an algorithmic breakthrough that shows how the underlying problem can be solved efficiently not only on a CRQC but on a classical computer as well. Therefore, as envisioned by the European Union Agency For Cybersecurity (ENISA) in its recent report on PQC (see details in (ENISA, 2022)), the cryptanalysts may have overlooked severe attacks on post-quantum candidates (digital signatures or key encapsulation mechanisms) that could be quickly conducted using today’s classical computers. In view of this, simply updating global networking protocols such as DNSSEC on a large scale with post-quantum digital signatures algorithms may be problematic because any attacks on these post-quantum candidates revealed in the near future will render the DNSSEC-based address resolution process untrustworthy. Conversely, keeping the conventional digital signature algorithms intact in DNSSEC (and not integrating post-quantum digital signatures) can also be problematic because CRQC inception may not be publicly announced (especially when any state or a nefarious actor builds CRQC sooner than anticipated). In such a situation, the sole reliance on conventional digital signatures for DNSSEC will make data integrity and source authentication unattainable. Therefore, the quantum threat as envisioned above makes a challenging and complex situation as of today - i.e., either sticking only to conventional digital signatures or switching only to post-quantum digital signatures can potentially have severe security consequences.

### v2

@@ Abstract.
DNSSEC, a DNS security extension, is essential to accurately translating domain names to IP addresses. Digital signatures provide the foundation for this reliable translation; however, the evolution of Quantum Computers has made traditional digital signatures vulnerable. In light of this, NIST has recently selected potential post-quantum digital signatures that can operate on conventional computers and resist attacks made with Quantum Computers. Since these post-quantum digital signatures are still in their early stages of development, replacing pre-quantum digital signature schemes in DNSSEC with post-quantum candidates is risky until the post-quantum candidates have undergone a thorough security analysis. Given this, herein, we investigate the viability of employing Double-Signatures in DNSSEC, combining a post-quantum digital signature and a classic one  [^1] A preliminary version of this article was accepted in ITNAC 2025. The rationale is that double-signatures will offer protection against quantum threats on conventional signature schemes as well as unknown non-quantum attacks on post-quantum signature schemes, hence even if one fails, the other provides security guarantees. However, the inclusion of two signatures in the DNSSEC response message doesn’t bode well with the maximum allowed size of DNSSEC responses (i.e., 1232B, a limitation enforced by the MTU of physical links). To counter this issue, we leverage a way to do application-layer fragmentation of DNSSEC responses with two signatures. We implement our solution on top of OQS-BIND and, through experiments, show that the addition of two signatures in DNSSEC and application-layer fragmentation of all relevant resource records and their reassembly does not have a substantial impact on the efficiency of the resolution process and thus is suitable for the interim period at least until the quantum computers are fully realized.
@@
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Conference; Nov 2025; Australiaisbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
The Domain Name System (DNS) is responsible for translating human-readable domain names into machine-understandable IP addresses. As can be imagined, any vulnerability in this system can potentially make such translation precarious, leading users to malicious servers instead of their intended destinations. The Security Extension of DNS - i.e., Domain Name System Security Extensions (DNSSEC) helps establish that the messages received by a client are from an authorized DNS server and have not been altered in transit. DNSSEC primarily leverages digital signatures for establishing the aforementioned properties of authenticity and integrity. However, the underlying security assumptions of conventional digital signature algorithms - i.e., the computational difficulty of integer factorization and discrete logarithms will not hold when an attacker has access to a Cryptographically Relevant Quantum Computer (CRQC) (Agency, 2021; Shor, 1994). To mitigate this problem, the National Institute of Standards and Technology (NIST) has recently selected three post-quantum digital signatures - i.e., FALCON (Prest et al., 2020), CRYSTALS-DILITHIUM (Lyubashevsky et al., 2020), and SPHINCS+ (Hulsing et al., 2020). An important envisioned property of these post-quantum digital signature algorithms is that they can run on classical computers but also withstand the attacks conducted using CRQCs, thereby offering sufficient protection against the impending threat of CRQCs. For DNSSEC to continue offering the needed authenticity and integrity in the post-quantum era, it must be updated with post-quantum digital signatures. However, for successful transitioning to the post-quantum era, the Interim Period - i.e., from now until CRQCs are fully realized or until the security of post-quantum candidates is fully established- presents a particular challenge. Precisely, in preparation for the post-quantum era, simply switching pre-quantum cryptography (such as conventional digital signatures in DNSSEC) with post-quantum cryptography (such as the post-quantum candidates selected by NIST) in the interim period may lower overall security (ENISA, 2022). The reason is that the selected post-quantum candidates (both key encapsulation mechanisms and digital signatures) have not yet gone through a thorough cryptanalysis against both classical and quantum attacks. The selected post-quantum candidates may be compromised even by conventional classical computers in the near future. Therefore, there is an inherent risk in simply switching from pre-quantum cryptography to post-quantum cryptography alone in that overall security may be lowered not only against CRQCs but also against today’s classical computers (see further details in Section 3).
[…]
@@
To circumvent the aforementioned issues, a potential alternative is to perform the fragmentation at the application layer (i.e., within the DNS protocol itself). While there exist ways to tackle the application layer fragmentation (see (Goertzen and Stebila, 2022; Rawat and Jhanwar, 2024)), none of them can accommodate double signatures and public keys. Therefore, to accomplish this, first, we constructed a Docker-based testbed on an Amazon EC2 instance using commercial-grade DNS software - i.e., BIND9 (OQS-BIND, see full details on our testbed in Section 5). Second, we explored generating two signatures (pre-quantum and post-quantum) on the Name Server using the default BIND9 tools. Third, we devised a way to fragment the DNSSEC responses with two signatures and public keys on the name servers and accurately reassemble them on the resolver. We also modified the source code of the BIND9 resolver to enable the verification of both signatures before marking the response as authenticated. Finally, using our testbed, we conducted an empirical analysis to show that our implementation can handle two signatures in DNSSEC simultaneously and quantify the impact of double signatures on the DNSSEC resolution process. Our experiments show that double signatures have a negligible impact on the average resolution time of DNSSEC compared to only post-quantum digital signatures. Therefore, this analysis confirms the efficacy of double signatures in DNSSEC for the interim period until CRQCs are fully realized. Our main contributions are summarized as follows:
@@
  • We thoroughly investigated the source code for BIND9 and modified the resolver component to enable the re-assembly and verifications of both pre-quantum and post-quantum digital signatures simultaneously.
  • Through empirical analysis, we show that the addition of double signatures in the DNSSEC resolution process did not affect the performance. Hence, it is recommended for the interim period until CRQCs are fully realized.
@@ 2. Domain Name System (DNS)
For an abstract-level understanding of the DNS translation process, imagine a client who queries the IP address of a domain name example.com. For this translation, the client will first send a query to a (caching) resolver which will handle all the subsequent translation processes on the client’s behalf. Assuming that the resolver has not previously cached this answer (i.e., the IP address of example.com), it will query the root server for the name server of .com domain names. Once the resolver receives the response, it queries this name server for the name server(s) of example.com. Finally, once the resolver knows the corresponding name server, it queries it for the IP address of the domain name example.com and returns it to the client.
As specified in RFC 1035, the generic DNS messages exchanged between the resolver and name servers have five sections (Mockapetris, 1987). Figure 1 shows an overview of a generic DNS message. The header section is always present and its structure is depicted in Figure 2. The Question section encapsulates the resolver’s query whose IP address needs to be resolved. This section has three main fields - i.e., QName (contains encoded query), QType (mentions the type of needed resource recorded in response), and QClass (contains the class of query, mostly specifies IN). As shown in Figure 1, the last three sections have a similar format - i.e., concatenation of various RRs. The RRs are actual data entries in the DNS database that provide information related to the queried domain name and help the actual translation of domain names to corresponding IP addresses. The top-level format of RRs is shown below in Figure 3. Going back to our earlier example of example.com, an RR with Name ‘example.com’ and TYPE ‘A’ will contain an IP address of 32 bits in the RDATA field, and with TYPE ‘NS’, the RDATA field will contain the domain name of the authoritative server for example.com. The Answer section in a response message contains RRs that answer the query from the resolver, while the Authority section contains RRs that point to the authoritative server. The Additional section contains RRs that somehow relate to the query but may not essentially contain the answers to the query.
@@
Figure 3. $R$’s Top-Level Format
@@
As explained above, the client initiates the translation process by querying the root server, then its children, and then its children’s children, and finally reaching the name server that provides the queried IP address. This hierarchical process facilitates an important element of DNSSEC - i.e., the chain of trust that helps establish trust in public keys needed for signature verification. For the chain of trust, each zone that is queried as part of the resolution process must have a digest of public KSK kept as the Delegation Signer resource record in its parent zone; otherwise, the public ZSK sent by the name server cannot be trusted. Since the root server(s) have no parent zone, it does not publish any delegation signer (DS) resource record. Therefore, for the entire chain of trust, the public KSK of the root server is obtained out-of-band so that the root server’s public KSK is pre-installed on most operating systems to eliminate the need for fetching and configuring the key(s). An overview of the chain of trust is shown in Figure 4.
@@
Since RRSIG and DNSKEY RRs are the root cause of the maximum message size issue related to the double signatures (see discussion in the introduction), below we present an overview of RRSIG and DNSKEY RRs (see Figures 5, 6). Their thorough understanding is pivotal for the simultaneous fragmentation of two signatures at the application layer. We refer the readers to RFC 4034 for a more detailed discussion on these relevant RRs (Arends et al., 2005).
Figure 5. RRSIG $R$ Wire Format
Figure 6. DNSKEY $R$ Wire Format
@@ 3.1. Quantum Threat
The assurances of source authenticity and data integrity in DNSSEC are underpinned by digital signatures. However, these conventional digital signatures used in DNSSEC rely on the complexity and difficulty of computational problems (e.g., integer factorization and discrete logarithms), which are difficult for attackers to solve using today’s classical computers (Nosouhi et al., 2023a, b). There exists strong evidence that these “computationally hard problems” will become solvable in “polynomial time” with the quantum computational model using the so-called CRQCs (Mosca, 2018; Agency, 2021). These CRQCs will facilitate the implementation of Shor’s Algorithm that will help solve the aforementioned computational problems in polynomial time (Shor, 1994), thereby rendering the conventional digital signatures insecure. For DNSSEC, it means that in the post-quantum era (i.e., when CRQCs will be fully realized), it would be difficult to establish the integrity and authenticity of all resource records exchanged as a part of the address resolution process. Therefore, in the post-quantum era, it will become impossible to trust the final IP address of a queried domain, as the conventional digital signature will no longer offer source authenticity and data integrity at an adequate level.
@@
Given the aforementioned impending issues with conventional digital signatures (and also with existing key encapsulation mechanisms), NIST is currently undertaking a Post Quantum Cryptography (PQC) standardization project. It aims to standardize digital signature algorithms and key encapsulation mechanisms that can run on today’s classical computers but can withstand the attacks conducted using the CRQCs (if they are realized in the future). In this pursuit, NIST has recently selected three post-quantum digital signature algorithms, including FALCON (Prest et al., 2020), CRYSTALS-DILITHIUM (Lyubashevsky et al., 2020), and SPHINCS++ (Hulsing et al., 2020). To make DNSSEC (or in general any other internet protocol) quantum-safe, it must be updated with the aforementioned post-quantum digital signature algorithms and be investigated for any issues that demand protocol-level changes to DNSSEC for the post-quantum era. However, switching from conventional digital signatures to post-quantum algorithms may be precarious. The reason is that the post-quantum digital signature algorithms as selected by NIST are currently in their infancy and thus have been studied and analyzed for a shorter period (as compared with their classical counterparts). First, there could be flaws in the algorithm design itself that make them vulnerable to cryptanalytic attacks. Second, side-channels in the different implementations could leak keys via timing attacks, power analysis attacks, and other similar attacks. Third, there could be an algorithmic breakthrough showing how the underlying problem can be solved efficiently on a CRQC and on a classical computer. Therefore, as envisioned by the European Union Agency for Cybersecurity (ENISA) in its recent report on PQC (see details in (ENISA, 2022)), the cryptanalysts may have overlooked severe attacks on post-quantum candidates (digital signatures or key encapsulation mechanisms) that could be quickly conducted using today’s classical computers. Given this, simply updating global networking protocols such as DNSSEC on a large scale with post-quantum digital signature algorithms may be problematic because any attacks on these post-quantum candidates revealed in the near future will render the DNSSEC-based address resolution process untrustworthy. Conversely, keeping the conventional digital signature algorithms intact in DNSSEC (and not integrating post-quantum digital signatures) can also be problematic because CRQC inception may not be publicly announced (especially when any state or a nefarious actor builds CRQC sooner than anticipated). In such a situation, the sole reliance on conventional digital signatures for DNSSEC will make data integrity and source authentication unattainable. Therefore, the quantum threat as envisioned above makes a challenging and complex situation as of today - i.e., either sticking only to conventional digital signatures or switching only to post-quantum digital signatures can potentially have severe security consequences.